OpenStudy (anonymous):

10. Determine the number of real solutions of -2x2 + 5x - 3 = 0. A. 0 B. 1 (double root) C. 2

OpenStudy (solomonzelman):

you can't factor this, completing the square would be hard to do; so I suggest to use a QUADRATIC FORMULA.

OpenStudy (anonymous):

okay and how would i set it up?

OpenStudy (solomonzelman):

\(\Large\color{blue}{ \bf \color{red} { a }x^2+\color{green} { b }x+\color{black} { c }=0 }\) \(\Large\color{blue}{ \bf \color{red} { -2 }x^2\color{green} { +5 }x+\color{black} { -3 }=0 }\) \(\Huge\color{blue}{ \bf \frac{-b± \sqrt{ b^2-4ac} }{2a} }\)

OpenStudy (solomonzelman):

See how to compare the "a", "b", "c" and how to plug them in ?

OpenStudy (solomonzelman):

The second equation is just -3, not +-3. sorry about that.

OpenStudy (anonymous):

im goingto try to solve it now

OpenStudy (solomonzelman):

yes, and I would prefer you to show the steps in here, if you will. you don't have to though.

OpenStudy (anonymous):

\[-3\sqrt{-2}^{2}-4(-2*3) \] @SolomonZelman

OpenStudy (solomonzelman):

you can use \frac{ }{ } as a fraction ♂ ♂ top bottom

OpenStudy (solomonzelman):

\(\Huge\color{purple}{ \sf ∠ \frac{-5±\sqrt{(5)^{2}- 4(-2)(-3) } }{2(-2)} }\)